OpenStudy (anonymous):

(5) (5) (5) does anybody know this if so please tell me

OpenStudy (anonymous):

125

OpenStudy (anonymous):

i have a pic if it would help

OpenStudy (the_fizicx99):

Like does it mean something or is it 5*5*5

OpenStudy (anonymous):

yes kind of like that but but what *'s 5?

OpenStudy (anonymous):

Well 5^3 is the same as (5) (5) (5), if that's what you mean.

OpenStudy (the_fizicx99):

" * " means times. Like x mean in 4 x 4

OpenStudy (anonymous):

5*5*5=5^3=125